OpenStudy (anonymous):

you roll a dice. what is the probability that you roll either a 5 or a 6?

OpenStudy (crashonce):

as there are 6 options rolling 5/6 is two choices which is 2/6 = 1/3

OpenStudy (anonymous):

umm..can you explain a bit more

OpenStudy (crashonce):

a dice has 6 sides

OpenStudy (crashonce):

5 or 6 means one of TWO sides

OpenStudy (crashonce):

so if there are 6 sides and you can roll 2

OpenStudy (crashonce):

that is 2 in 6 chance or 1 in 3 = 1/3

OpenStudy (crashonce):

is that clear?

sammixboo (sammixboo):

If you roll a die you would have a \(\frac{1}{6}\) of a chance of rolling a 5 and same for rolling a 6, so therefore you would have \(\frac{2}{6}\) of a chance of rolling a 5 or a 6

Miracrown (miracrown):

You can't roll a ''die'' - sammixboo ;)

sammixboo (sammixboo):

Or a \(\frac{1}{3}\) of a chance, because \(\frac{2}{6}\) simplified is \(\frac{1}{3}\_

sammixboo (sammixboo):

Shh I know

sammixboo (sammixboo):

Dang it a LaTeX fail I must be getting tired

OpenStudy (uri):

LOL

sammixboo (sammixboo):

Uri shhhh

sammixboo (sammixboo):

But anyways \(\frac{2}{6}\) is simplified to \(\frac{1}{3}\)

OpenStudy (anonymous):

i get it know

sammixboo (sammixboo):

Okie Dokie :) if you need any more help you can tag me or anyone else by typing @ then their username Example : @jackelyn_1234

OpenStudy (uri):

iz okay gurl

OpenStudy (anonymous):

You have tiles and spell out the word ANACONDA. you scramble them and pick one at random. What is the chance that you pick an A? i believe it is 3/8 jut wanted to see if i'm correct

sammixboo (sammixboo):

Hehe, that word? Weird. Anyways! ANACONDA has 8 letters, and there is 3 'A's in that word, so therefore you would have a \(\frac{3}{8}\) chance of getting an 'A'

sammixboo (sammixboo):

Or a 37.5% chance :P

OpenStudy (anonymous):

so i was correct. now this one is really trowing me off. You have 3 card deck containing a King , Queen,and a Jack. you pick a card AND PUT IT BACK and pick another one. What is the probability that you do not pick the same card?

sammixboo (sammixboo):

Well if you have 3 cards, draw one, and then put it back you will have the same amount of cards. Let's say you chose a King on your first draw. Out is 3 cards there is one king, so when you draw a second time you would have a \(\frac{1}{3}\) of a chance to get a king making it a \(\frac{2}{3}\) of a chance of drawing another card

OpenStudy (anonymous):

so the chances of not picking the same card is 2/3?